# A theory of quasiballistic spin transport

### v2

@@ IV Discussion
We have developed a theory of quasiballistic spin transport, showing that the analytical approach proposed for charge transport in Ref. [16] can be extended to spin transport and qualitatively explains the numerical findings of Refs. [14, 15]. Our analysis of this phenomenon in the perturbative regime of large $\alpha$ further suggests that the observed quasiballistic behaviour is enabled by a non-trivial interplay between hopping and interaction terms that suppresses the decay of the spin current more strongly than one might naïvely expect.
@@
This analytical approach is theoretically justified for the power-law interacting XXZ models in this paper provided they are not “secretly” ballistic away from the known integrable points $\alpha=2$ and $\alpha\to\infty$. More precisely, our approach rules out perfectly ballistic spin transport (i.e. exact conservation of the spin current operator) away from these points, but it does not rule out imperfectly ballistic spin transport, as exemplified by the nearest-neighbour spin-$1/2$ XXZ chain with anisotropy $10<|\Delta|<$ 1. In the latter case, the spin current is not exactly conserved, but it has sufficiently large overlap with a conserved operator in the thermodynamic limit that the Mazur bound [34, 35] guarantees ballistic spin transport [32, 33] and the approach taken in this paper becomes invalid.
@@
While we cannot exclude this possibility based on available theoretical techniques, we prove a partial result justifying our approach in Appendix A.2. Specifically, we show that for the models studied in this paper, there is no two-body conserved charge that could give rise to a non-zero Mazur bound for $\Delta\neq 0,\pm 1$. Unfortunately, this does not eliminate the possibility of quasilocal, many-body conserved charges that protect the spin current, as in the nearest-neighbour XXZ chain [33, 36]. At the same time, a systematic construction of quasilocal or pseudolocal [36] conserved charges for an arbitrary lattice model without prior knowledge of its underlying integrable structure is an extremely challenging problem whose solution would have immediate implications for various important open problems in many-body physics, such as the rigorous justification of many-body localization [37, 38, 39, 40].
@@
Meanwhile, there is a long-standing expectation in the literature that ballistic transport of some local conserved charge should always follow from integrability; this belief is sometimes referred to as the “integrability-transport conjecture” [7, 32, 4]. While ballistic energy transport in the nearest-neighbour XXZ model is consistent with this conjecture, the precise relationship between integrability and spin transport is murkier [6]. We propose the following revision of the integrability-transport conjecture for inversion-symmetric lattices of spins: spin transport is ballistic at all temperatures only if the system is integrable. We do not know of any counterexamples to this conjecture, but it seems difficult to prove for the reasons discussed above.
@@
A corollary of our analysis is a simplified integrability-transport theorem (see Theorem 1 of Appendix A.1): for spin-$1/2$ chains of the form Eq. (1), we show that spin transport is only perfectly ballistic if the model is equivalent to either a nearest-neighbour XX chain or a Haldane-Shastry chain. Extending this result to imperfect ballistic transport and pseudolocal conserved charges, along the lines of recent rigorous results on non-integrability of certain spin chains [41, 42, 43], seems desirable for understanding the relationship between integrability and transport more generally.
@@ A.1 Necessary and sufficient condition for perfectly ballistic spin transport
$(u-v)a(u)a(v)+(u+v)a(u+v)(b(u)-b(v))=0$ (15)
holds for all $u,v\in\mathbb{Z}$ such that $u,\,v,\,u+v\neq 0$. This functional equation is satisfied by both the nearest-neighbour XX and Haldane-Shastry models of interest in this work, which therefore exhibit perfectly ballistic spin transport. We now show that provided $b(r)\to 0$ as $r\to\infty$ (which is necessary on physical grounds) these are the only solutions to Eq. (15), up to a choice of sublattice scale and a standard [1] spin-flip-type unitary transformation on even-parity sites of each sublattice.
@@
Next suppose that $b$ is non-zero, so there exists some $x>0$ such that $b(x)\neq 0$. Suppose for a contradiction that the support of $a$ is bounded, i.e. $N_{a}=\max(E_{a})<\infty$. Letting $u=x+N_{a}$, $v=-x$ in Eq. (15), it follows that $x+N_{a}\notin E_{a}$ so that
@@
We now show that $E_{b}\subseteq E_{a}$. Suppose there exists some $u\in E_{b}\setminus E_{a}$. Then $u=mR+k$ for integers $m\geq 0$ and $0<k<R$. Let $v=nR-k$ for some integer $n>m$. Then $u,v\notin E_{a}$ but $u+v=(m+n)R\in E_{a}$. Substituting into Eq. (15), we deduce that $b(mR+k)=b(nR-k)\neq 0$ for all integers $n>m$, which contradicts the decay of $b(r)$. Thus $E_{b}\subseteq E_{a}$.
It remains to determine the specific behaviour of $a(r)$ and $b(r)$ on $E_{a}$. To this end, let $u=nR$ and $v=\pm R$ in Eq. (15) for any integer $n>1$. Then
@@
Then, picking any $l>m>n$, the identity $b((2l+1)R)-b((2n+1)R)=(b((2l+1)R)-b(2mR))+(b(2mR)-b((2n+1)R))$ demands that $\lambda^{2}=1$.
First consider the case $\lambda=1$. It follows by the above equations that $b(mR)-b(nR)=a(mR)-a(nR)$ for all positive integers $m$ and $n$. In particular,
@@ A.2 Sufficient condition for a non-zero Mazur bound
To justify the theoretical approach taken in the main text, we would ideally like to rule out the existence of such conserved operators rigorously. In this Appendix, we make partial progress along these lines and show that for the power-law interacting XXZ chains studied in this paper, there is no two-body conservation law $\hat{K}$ that could yield a non-zero Mazur bound for spin transport for model anisotropies $\Delta\neq 0,\pm 1$. While the restriction to two-body operators does not capture the phenomenology of e.g. the nearest-neighbour XXZ chain [33], it does capture the phenomenology of the Haldane-Shastry and nearest-neighbour XX models relevant to this work.
@@
where $K$ is an odd function of its argument $K(-r)=-K(r)$. Requiring that $\hat{K}$ be a conserved operator leads to a nontrivial functional equation connecting $a(r)$, $b(r)$, and $K(r)$, namely
$K(u)a(v)-K(v)a(u)+K(u+v)(b(u)-b(v))=0$ (33)
@@
Multiplying this equation by $a(x)a(y)(b(x)-b(y))$ finally yields the functional equation

### v3

@@ IV Discussion
We have developed a theory of quasiballistic spin transport, showing that the analytical approach proposed for charge transport in Ref. [HBS] can be extended to spin transport and qualitatively explains the numerical findings of Refs. [Mierzejewski_2023, Yang_2024]. Our analysis of this phenomenon in the perturbative regime of large $\alpha$ further suggests that the observed quasiballistic behavior is enabled by a non-trivial interplay between hopping and interaction terms that suppresses the decay of the spin current more strongly than one might naïvely expect.
@@
This analytical approach is reasonable for the power-law interacting XXZ models in this paper provided they are not “secretly” ballistic away from the known integrable points $\alpha=2$ and $\alpha\to\infty$. More precisely, our approach rules out perfectly ballistic spin transport (i.e. exact conservation of the spin current operator) away from these points, but it does not rule out imperfectly ballistic spin transport, as exemplified by the nearest-neighbour spin-$1/2$ XXZ chain with anisotropy $0<|\Delta|<1$. In the latter case, the spin current is not exactly conserved, but it has sufficiently large overlap with a conserved operator in the thermodynamic limit that the Mazur bound [mazur1969non, suzuki1971ergodicity] guarantees ballistic spin transport [Zotos_1997, ProsenBound] and the approach taken in this paper becomes invalid.
@@
While we cannot exclude this possibility based on available theoretical techniques, we prove a partial result justifying our approach in Appendix A.2. Specifically, we show that for the models studied in this paper, there is no two-body conserved charge that could give rise to a non-zero Mazur bound for $\Delta\neq 0,\pm 1$. Unfortunately, this does not eliminate the possibility of quasilocal, many-body conserved charges that protect the spin current, as in the nearest-neighbour XXZ chain [ProsenBound, Ilievski_2016]. At the same time, a systematic construction of quasilocal or pseudolocal [Ilievski_2016] conserved charges for an arbitrary lattice model without prior knowledge of its underlying integrable structure is an extremely challenging problem whose solution would have immediate implications for various important open problems in many-body physics, such as the rigorous justification of many-body localization [Basko_2006, OganesyanHuse, VoskAltman, deroeck2023rigoroussimpleresultsslow].
@@
Meanwhile, there is a long-standing expectation in the literature that ballistic transport of some local conserved charge should always follow from integrability; this belief is sometimes referred to as the “integrability-transport conjecture” [Castella_1995, Zotos_1997, Sirker_2011]. While ballistic energy transport in the nearest-neighbour XXZ model is consistent with this conjecture, the precise relationship between integrability and spin transport is murkier [Bertini]. We propose the following revision of the integrability-transport conjecture for inversion-symmetric lattices of spins: spin transport is ballistic at all temperatures only if the system is integrable. We do not know of any counterexamples to this conjecture, but it seems difficult to prove for the reasons discussed above.
@@
A corollary of our analysis is a simplified integrability-transport theorem (see Theorem 1 of Appendix A.1): for spin-$1/2$ chains of the form Eq. (1), we show that spin transport is only perfectly ballistic if the model is equivalent to either a nearest-neighbour XX chain or a Haldane-Shastry chain. Extending this result to imperfect ballistic transport and pseudolocal conserved charges, along the lines of recent rigorous results on non-integrability of certain spin chains [yamaguchi2024completeclassificationintegrabilitynonintegrability, yamaguchi2024proofabsencelocalconserved, shiraishi2025complete], seems desirable for understanding the relationship between integrability and transport more generally.
@@ A.1 Necessary and sufficient condition for perfectly ballistic spin transport
(uv)a(u)a(v)+(u+v)a(u+v)(b(u)b(v))=0(u-v)a(u)a(v)+(u+v)a(u+v)(b(u)-b(v))=0 (15)
holds for all $u,v\in\mathbb{Z}$ such that $u,\,v,\,u+v\neq 0$. This functional equation is satisfied by both the nearest-neighbour XX and Haldane-Shastry models of interest in this work, which therefore exhibit perfectly ballistic spin transport. We now show that provided $b(r)\to 0$ as $r\to\infty$ (which is necessary on physical grounds) these are the only solutions to Eq. (15), up to a choice of sublattice scale and a standard [takahashi2005thermodynamics] spin-flip-type unitary transformation on even-parity sites of each sublattice.
@@
Next suppose that $b$ is non-zero, so there exists some $x>0$ such that $b(x)\neq 0$. Suppose for a contradiction that the support of $a$ is bounded, i.e. $N_{a}=\max(E_{a})<\infty$. Letting $u=x+N_{a}$, v=xv=-x in Eq. (15), it follows that $x+N_{a}\notin E_{a}$ so that
@@
We now show that $E_{b}\subseteq E_{a}$. Suppose there exists some $u\in E_{b}\setminus E_{a}$. Then $u=mR+k$ for integers $m\geq 0$ and $0<k<R$. Let v=nRkv=nR-k for some integer $n>m$. Then $u,v\notin E_{a}$ but $u+v=(m+n)R\in E_{a}$. Substituting into Eq. (15), we deduce that $b(mR+k)=b(nR-k)\neq 0$ for all integers $n>m$, which contradicts the decay of $b(r)$. Thus $E_{b}\subseteq E_{a}$.
It remains to determine the specific behavior of $a(r)$ and $b(r)$ on $E_{a}$. To this end, let $u=nR$ and $v=\pm R$ in Eq. (15) for any integer $n>1$. Then
@@
Then, picking any $l>m>n$, the identity b((2l+1)R)b((2n+1)R)=(b((2l+1)R)b(2mR))+(b(2mR)b((2n+1)R))b((2l+1)R)-b((2n+1)R)=(b((2l+1)R)-b(2mR))+(b(2mR)-b((2n+1)R)) demands that $\lambda^{2}=1$.
First consider the case $\lambda=1$. It follows by the above equations that b(mR)b(nR)=a(mR)a(nR)b(mR)-b(nR)=a(mR)-a(nR) for all positive integers $m$ and $n$. In particular,
@@ A.2 Sufficient condition for a non-zero Mazur bound
To justify the theoretical approach taken in the main text, we would ideally like to rule out the existence of such conserved operators rigorously. In this Appendix, we make partial progress along these lines and show that for the power-law interacting XXZ chains studied in this paper, there is no two-body conservation law $\hat{K}$ that could yield a non-zero Mazur bound for spin transport for model anisotropies $\Delta\neq 0,\pm 1$. While the restriction to two-body operators does not capture the phenomenology of e.g. the nearest-neighbour XXZ chain [ProsenBound], it does capture the phenomenology of the Haldane-Shastry and nearest-neighbour XX models relevant to this work.
@@
where $K$ is an odd function of its argument K(r)=K(r)K(-r)=-K(r). Requiring that $\hat{K}$ be a conserved operator leads to a nontrivial functional equation connecting $a(r)$, $b(r)$, and $K(r)$, namely
K(u)a(v)K(v)a(u)+K(u+v)(b(u)b(v))=0K(u)a(v)-K(v)a(u)+K(u+v)(b(u)-b(v))=0 (33)
@@
Multiplying this equation by a(x)a(y)(b(x)b(y))a(x)a(y)(b(x)-b(y)) finally yields the functional equation